LASER-INDUCED FLUORESCENCE SPECTROSCOPY OF THE Ã ^{2}Π − X̃ ^{2}Σ^{+} TRANSITION OF SrCCH AND SrCCD

A. FARIDIAN, \underline{D. W. TOKARYK}, Department of Physics and Centre for Laser, Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, Canada E3B 5A3; A. G. ADAM, Chemistry Department, and Centre for Lasers, and Atomic, and Molecular Sciences, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

The family of molecules consisting of an alkaline earth metal atom bonded to an organic ligand has been under intensive investigation recently, since these molecules play a significant role in synthesis of organic compounds, and since some members like MgNC and MgCN have been detected in stellar atmospheres. Our group has previously studied CaCCH, MgCCH, CaOCH_{3}, and SrOCH_{3}, and in this work we consider the Ã ^{2}Π − X̃ ^{2}Σ^{+} transition of SrCCH and SrCCD. The molecules were created in a pulsed-jet laser ablation source through reaction of strontium atoms ablated from a solid rod with methane gas entrained in helium. High-resolution spectra were collected through laser-induced fluorescence following excitation of the molecules formed in the jet with a cw ring dye laser. Results of the analysis are compared to those of a previous high-resolution study of SrCCH.^{a} The data provide information on the C-H bond length via the isotopic substitution, and the lambda-doubling parameter p allows for an estimate of the position of the B̃ ^{2}Σ^{+} state for both isotopes.

aM. J. Dick, P. M. Sheridan, J.-G. Wang, and P. F. Bernath , J. Mol. Spectrosc. \underline{233}, pp. 197-202 (2005)